# Bit of a Close Talker: A Practical Guide to Serverless Cloud Co-Location Attacks

### v1

@@ Abstract
Serverless computing has revolutionized cloud computing by offering an efficient and cost-effective way for users to develop and deploy applications without managing infrastructure details. However, serverless cloud users remain vulnerable to various types of attacks, including micro-architectural side-channel attacks. These attacks typically rely on the physical co-location of victim and attacker instances, and attackers will need to exploit cloud schedulers to achieve co-location with victims. Therefore, it is crucial to study vulnerabilities in serverless cloud schedulers and assess the security of different serverless scheduling algorithms. This study addresses the gap in understanding and constructing co-location attacks in serverless clouds. We present a comprehensive methodology to uncover exploitable features in serverless scheduling algorithms and devise strategies for constructing co-location attacks through normal user interfaces. In our experiments, we successfully reveal exploitable vulnerabilities and achieve instance co-location on prevalent open-source infrastructures and Microsoft Azure Functions. We also present a mitigation strategy to defend against co-location attacks in serverless clouds. Our work highlights critical areas for security enhancements in current cloud schedulers, offering insights to fortify serverless computing environments against potential co-location attacks.
@@ I Introduction
In recent years, serverless computing, a relatively new form of cloud computing, has become prevalent and attracts a great number of users. It fully offloads the task of infrastructure management to cloud providers, enabling developers to focus more on the functionality of their applications and to host services in a more cost-efficient way [42, 27]. Today, we can still see an increasing number of commercial services pivoting to this cloud computing model. This computing paradigm has been supported by mainstream cloud providers, e.g. Amazon AWS Lambda [6], Google Cloud Run [20], and Microsoft Azure Functions [31]. Compared to traditional cloud computing (often referred to as “serverful computing”), serverless computing has the following benefits: (1) Flexibility, as the cloud provider-managed light-weight serverless instances can flexibly scale depending on the load of clients’ services and cloud providers can achieve higher resource efficiency; (2) Convenience, since client developers now only need to develop functional code and directly submit code or container images to cloud providers, with all resource management details being hidden to them; (3) Customer cost efficiency, as customers are only charged for the actual resources they utilize (often measured by the number of function invocations), unlike in serverful cloud computing where customers are charged by instance time and a lot of costs are wasted on instance idling.
@@
Resource provisioning, i.e. scheduling, is an important aspect of serverless cloud operation. It impacts the performance of customer applications and can harm the profits of cloud providers if serverless instances are not properly placed and provisioned. Apart from resource scheduling challenges that are similar to serverful clouds, one unique challenge in serverless scheduling is the existence of cold-start [34], i.e. the time-consuming start-up process of serverless instances. Another unique challenge arises from the flexible nature of resource management of serverless systems, and serverless cloud providers need to adaptively perform auto-scaling to accommodate the dynamic service workloads. There are lots of research works dedicated to improving the performance of serverless systems from scheduling policies [19, 28, 4].
@@
Despite being convenient and cost-effective for cloud users, serverless clouds still face security challenges posed by various types of micro-architectural side-channel attacks. For example, remote micro-architectural attacks like Prime+Probe [15] remain feasible in serverless clouds with certain optimizations [48], threatening the security of serverless users. These attacks depend on the co-location of attackers’ and victims’ serverless instances on the same physical server within the cloud cluster. Although perfect isolation can be obtained by switching to dedicated hosts, this comes at a prohibitive cost: the cheapest dedicated servers from AWS and Azure start at \$323 and \$444 per month respectively, which is over 10 times higher than running the same workload on serverless platforms, as discussed in Section IX-B. Therefore, beyond understanding the mechanics of the actual attack, it is crucial to study how this critical prerequisite of co-location can be achieved and to analyze potential attacker strategies. This analysis will inspire the design of more secure serverless schedulers. Co-location attack methods have been investigated in traditional clouds [18, 30, 40, 44] and Google Cloud Run [47]. However, there is still a lack of studies on a universal methodology to exploit serverless cloud schedulers.
@@
In this work, we aim to bridge this gap in the literature and provide universal guidance on the construction of co-location attacks targeting different serverless clouds. We first provide a method to reveal exploitable features of serverless cloud scheduling algorithms, such as package locality optimizations [8]. Then, based on the found features, we offer a strategy to construct attacks accordingly. All these steps are done through a normal user interface, i.e. besides deploying serverless applications and invoking corresponding services, attackers do not have other access permissions to the cloud infrastructure. The main design effort is hence dedicated to constructing a set of attacks against serverless services from the user side.
@@
  • Additionally, we propose a scheduling algorithm called Double-Dip that aims to provide defence by offering “soft” user isolation.
@@ II-A Serverless Schedulers
Serverless scheduling is different from traditional scheduling in serverful clouds. In serverless clouds, user code is executed in encapsulated environments that pack required dependencies, e.g. containers or vendor-specific instances [2]. Serverless schedulers are responsible for deciding the placements of these instances. Like in traditional cloud scheduling and cluster scheduling, serverless schedulers need to determine how resources are assigned to user instances in the system based on resource availability. Additionally, serverless schedulers need to make decisions on how to dispatch serverless function invocations to instances. A major thread of research is dedicated to optimizing start time and avoiding unnecessary cold-start latency by optimizing scheduling policies [3, 43, 41]. Besides, as serverless cloud providers are also responsible for dynamically scaling function instances [33], auto-scaling and load-balancing are also important aspects of serverless schedulers. There are research works that consider different kinds of locality [1, 8] to optimize for load balancing performance. Later in this paper, we will show that from an attacker’s view, these are features that can be exploited to increase the co-location attack success rate.
@@ II-B Cloud Co-Location Attacks
Step 3 is the focus of most side-channel attack works, while Step 0 - Step 2 are usually neglected by researchers. However, achieving co-location is a non-trivial process, especially as the size of datacenters increases over time. Prior works in recent years [30, 18, 47] indicate that special strategies should be constructed according to scheduler features to tamper with targeted cloud systems.
@@ II-D Threat Model
In this paper, we only study Step 0 (Section III) - Step 1 (Section V) of the co-location attack. We assume non-privileged attackers, i.e. attackers can only launch attacks by invoking functions and collecting the returned execution results. The goal of attackers is to place at least one of their attack function hosts together with a victim’s function host on the same physical machine to enable subsequent attack efforts. In addition, we assume that the attacker does not have any knowledge about the targeted serverless scheduler. Scheduling strategies need to be reverse-engineered (using our method in Section III), based on which the attacker constructs an attack strategy accordingly to target a specific scheduler. However, they can have certain knowledge about the targeted application, e.g. dependencies, which is realistic in real world. Examples include Airbnb’s disclosure of their usage of open-source projects [5] and Netflix’s publication of their internal tool as an open-source project [36].
@@ III-A Algorithm Description
In this part, we assume a reliable server identification method is provided, i.e. when invoking functions, the attacker will be able to differentiate servers where the function host is placed. This can be based on methods such as time and frequency probing [47] or network-based probing [40]. With this server fingerprinting method, an attacker can continuously ‘discover’ servers in the cloud system and provide unique identifiers to each discovered physical server. Following the same function invocation sequence, the collected trace of such identifiers will be different for different scheduling algorithms, hence they can be used to identify scheduling policies utilized by the cloud provider.
We denote the aforementioned server identification method as a function $\mathcal{F}$. Every time a serverless function $\phi$ is executed in the cluster, a hypothetical fingerprint of the corresponding physical server will be reported [denoted as $\mathcal{F}(\phi)$]. We maintain a mapping record ($\mathcal{R}$) between physical server fingerprints and physical server IDs. When a fingerprint that does not map to any server is returned, a new server in the cluster is discovered. A new ID will be assigned to the server and a new entry in $\mathcal{R}$ will be established to record the mapping between the server fingerprint and the assigned server ID. The trace of server IDs associated with returned server fingerprints after each function invocation is recorded, and will later be used for analysis. As an example, consider a function that is invocated $6$ times. If the first $3$ executions are handled by one server in the cluster, and the last $3$ executions are on another server, the resulting sequence will be $(1,1,1,2,2,2)$.
The submission of function invocation consists of $3$ phases (a detailed pseudocode description can be found in Appendix A):
@@
Phase 2. In this step, the attacker constructs a second serverless function $\phi_{1}$, which is the copy of the previous serverless function $\phi_{0}$. Then, the attacker will invoke the two functions alternately for a specific length of time (e.g. 10 minutes), and then wait for all function hosts to be terminated.
@@
Analysis. Features of the scheduling algorithm can be revealed by analyzing and comparing different parts of the collected trace. In Phase 1, the attackers invoke the same function with a relatively high frequency for a sufficiently long pre-defined length of time. The placement sequence in the collected trace can reveal the following features ($\mathrm{F}_{i}$) of the scheduler:
@@
In Phase 2, after all function hosts are terminated, the attacker re-invokes the same function in Phase 1, together with another function with the same functionality but disguised as a different function. This enables the attacker to compare placement results between (1) invocations to the same function that are separated apart in time; and (2) invocations to different functions. The attacker will be able to answer the following questions:
@@
    Cold-Start Locations: By comparing the placement results of the first function in Phase 1 and Phase 2, the attacker will be able to obtain information regarding the scheduler’s behavior when re-scheduling a function instance that has been terminated before. The attacker will be able to answer the following question: for the same function, is the cold-start location always the same?
@@
    Account Locality: By comparing placement results of different functions, the attacker will be able to infer if user locality is considered by the scheduler and answer this question: do the scheduler tend to place function hosts from the same user on the same set of physical servers? This can potentially infect the construction of an attack strategy. This also needs to be cross-checked with placement results from Phase 3.
The goal of Phase 3 is to explore whether there are other locality optimizations that rely on specific configurations of serverless functions (e.g. package required) and are employed by the scheduler. By comparing the placement results of different functions, the attacker will be able to identify:
@@ IV-A Simulator Details
We conduct our fingerprinting experiments in a simulator written in C++. The usage of this simulator allows us to launch large-scale Monte Carlo experiments. The simulator consists of $\sim$2000 LoC and simulates the functionality of a serverless system, with a focus on the scheduler. There are $4$ different components in the simulator:
@@
At the beginning of the simulation, a list of users will be randomly generated and an attacker will be initiated to target a specific user. Both non-malicious users and attackers will generate Function instances when they are initialized. The simulation will go through a pre-defined number of rounds of Function submissions by different users, and the log files will be collected for further analysis. We have implemented different scheduling algorithms in different Scheduler modules.
@@ IV-B Selected Scheduling Algorithms
We select to implement the following schedulers in both the simulator and the Dask-based serverless platform:
@@
    Helper, which mimics the critical load-balancing behavior of Google Cloud Run [20] described in [47]. For each function to be scheduled, the scheduler selects a server as the base server at first. As the load increases, the scheduler creates function host instances on other servers to perform auto-scaling.
@@
    PASch [8], which is a serverless algorithm that utilizes package locality to improve the performance of serverless platforms. The scheduler maintains a consistent hash ring [24] during operation and maps serverless functions to servers based on the largest package that is required to increase package cache hit rate in the system and at the same time balance the load of each server in the system.
@@
We utilize OpenWhisk [37] as an example of cold-start-optimized schedulers. Also, as it is one of the most popular open-source serverless platforms, our attack results will be representative. We select Helper [47] since it is reported to be part of the scheduling algorithm of Google Cloud Run and we also observe similar behaviors in Microsoft Azure (see Section VII). PASch is selected to represent schedulers with $\mathrm{F}_{5}$ (configuration-based locality) optimizations, and the attack methods on PASch can be easily expanded to target other schedulers with similar optimization features. Finally, Random is selected as the baseline and helps demonstrate if increased randomness can help defend against co-location attacks, as shown by [17] in serverful clouds.
@@ IV-B1 Results
Helper Scheduler. By examining the placement trace in Phase 1, we can clearly observe that the scheduler has optimizations to reduce cold-starts, as functions are invoked repeatedly on the same set of servers ($\mathrm{F}_{1}$). As the number of function invocations increases, the scheduler starts to scale out function hosts by creating more instances to be placed on other servers ($\mathrm{F}_{2}$). In Phase 2, we can obtain the same observation as we start to involve a second function. By comparing the placement results of the overlapped function in Phase 1 and Phase 2, we find that if all hosts of a function are terminated, when re-launching these function hosts the target scheduler will still choose the same set of servers, i.e. the Helper scheduler presents cold-start locality ($\mathrm{F}_{3}$). Also, we notice that the two different functions in Phase 2 are placed in different locations, indicating the Helper function does not confine the host placement to a small set of servers even though they are from the same user ($\mathrm{F}_{4}$). However, the behaviors of the Helper scheduler in Phase 3 are similar to Phase 1 & 2, indicating that when making scheduling decisions the scheduler does not consider package locality ($\mathrm{F}_{5}$).
@@
OpenWhisk Scheduler. In Phase 1, when there is only one function involved, we can see that the function executions consistently stay on one machine, indicating that there is optimization to reduce cold-start ($\mathrm{F}_{1}$) yet no auto-scaling involved ($\mathrm{F}_{2}$). In Phase 2, as we involve a second function, we can see that the number of covered servers increases by 1, indicating that function hosts from the same user are not bonded together ($\mathrm{F}_{4}$), and the location is the same when a function host experiences termination and cold-start ($\mathrm{F}_{3}$). In Phase 3, however, we do not notice the scheduling results change, indicating that the OpenWhisk scheduler does not use package information when making scheduling decisions.
@@
PASch Scheduler. Similar to the OpenWhisk scheduler, in Phase 1, traces collected from PASch show instance locality ($\mathrm{F}_{1}$). However, since the placement consistently stays at the same machine, we can conclude that it does not perform auto-scaling based on the function invocation load ($\mathrm{F}_{2}$). In Phase 2, when a second function is involved, there is still only one server covered. This could potentially be due to (1) user locality, i.e. the scheduler places all functions from a user to a single machine; or (2) the identical configurations of the two functions result in the same placement decision. By examining Phase 3 of the trace, we can exclude possibility (1) since there are different function hosts in this phase. Besides, by comparing scheduling results in Phase 3, an attacker would be able to notice that the scheduler places function hosts with the same package requirements at the same node, demonstrating the existence of package locality-related policies in the scheduling algorithm.

### v3

@@ Abstract
Serverless computing has revolutionized cloud computing by offering users an efficient, cost-effective way to develop and deploy applications without managing infrastructure details. However, serverless cloud users remain vulnerable to various types of attacks, including micro-architectural side-channel attacks. These attacks typically rely on the physical co-location of victim and attacker instances, and attackers need to exploit cloud schedulers to achieve co-location with victims. Therefore, it is crucial to study vulnerabilities in serverless cloud schedulers and assess the security of different serverless scheduling algorithms. This study addresses the gap in understanding and constructing co-location attacks in serverless clouds. We present a comprehensive methodology to uncover exploitable features in serverless scheduling algorithms and to devise strategies for constructing co-location attacks via normal user interfaces. In our experiments, we successfully reveal exploitable vulnerabilities and achieve instance co-location on prevalent open-source infrastructures and Microsoft Azure Functions. We also present a mitigation strategy, the Double-Dip scheduler, to defend against co-location attacks in serverless clouds. Our work highlights critical areas for security enhancements in current cloud schedulers, offering insights to fortify serverless computing environments against potential co-location attacks.
@@ I Introduction
In recent years, serverless computing, a relatively new form of cloud computing, has become prevalent and attracts a great number of users. It fully offloads the task of infrastructure management to cloud providers, enabling developers to focus more on the functionality of their applications and to host services more cost-effectively [42, 27]. Today, we can still see an increasing number of commercial services pivoting to this cloud computing model. This computing paradigm has been supported by mainstream cloud providers, e.g., Amazon AWS Lambda [7], Google Cloud Run [20], and Microsoft Azure Functions [33]. Compared to traditional cloud computing (often referred to as “serverful computing”), serverless computing has the following benefits: (1) Flexibility, as the cloud provider-managed light-weight serverless instances can flexibly scale depending on the load of clients’ services and cloud providers can achieve higher resource efficiency; (2) Convenience, since client developers now only need to develop functional code and directly submit code or container images to cloud providers, with all resource management details being hidden from them; (3) Customer cost efficiency, as customers are only charged for the actual resources they utilize (often measured by the number of function invocations), unlike in serverful cloud computing, where customers are charged by instance time and a lot of costs are wasted on instance idling.
@@
Resource provisioning, i.e., scheduling, is an important aspect of serverless cloud operation. It impacts the performance of customer applications and can harm the profits of cloud providers if serverless instances are not properly placed and provisioned. Apart from resource scheduling challenges that are similar to serverful clouds, one unique challenge in serverless scheduling is the existence of cold-start [34], i.e., the time-consuming start-up process of serverless instances. Another unique challenge arises from the flexible nature of resource management of serverless systems, and serverless cloud providers need to adaptively perform auto-scaling to accommodate the dynamic service workloads. There are lots of research works dedicated to improving the performance of serverless systems from scheduling policies [19, 28, 4].
@@
Despite being convenient and cost-effective for cloud users, serverless clouds still face security challenges posed by various types of micro-architectural side-channel attacks. For example, remote micro-architectural attacks like Prime+Probe [15] remain feasible in serverless clouds with certain optimizations [48], threatening the security of serverless users. These attacks depend on the co-location of attackers’ and victims’ serverless instances on the same physical server within the cloud cluster. Although perfect isolation can be obtained by switching to dedicated hosts, this comes at a prohibitive cost: the cheapest dedicated servers from AWS and Azure start at \$323 and \$444 per month, respectively, which is over 10 times higher than running the same workload on serverless platforms, as discussed in Section IX-B. Therefore, beyond understanding the mechanics of the actual attack, it is crucial to study how this critical prerequisite of co-location can be achieved and to analyze potential attacker strategies. This analysis will inspire the design of more secure serverless schedulers. Co-location attack methods have been investigated in traditional clouds [18, 30, 40, 44] and Google Cloud Run [47]. However, there is still a lack of studies on a universal methodology to exploit serverless cloud schedulers.
@@
In this work, we aim to bridge this gap in the literature and provide universal guidance on the construction of co-location attacks targeting different serverless clouds. We first provide a method to reveal exploitable features of serverless cloud scheduling algorithms, such as package locality optimizations [8]. Then, based on the found features, we offer a strategy to construct attacks accordingly. All these steps are done through a normal user interface, i.e., besides deploying serverless applications and invoking corresponding services, attackers do not have other access permissions to the cloud infrastructure. The main design effort is hence dedicated to constructing a set of attacks against serverless services from the user side.
@@
  • Additionally, we propose a scheduling algorithm called Double-Dip that aims to provide defense by offering “soft” user isolation.
@@ II-A Serverless Schedulers
Serverless scheduling is different from traditional scheduling in serverful clouds. In serverless clouds, user code is executed in encapsulated environments that pack required dependencies, e.g., containers or vendor-specific instances [2]. Serverless schedulers are responsible for deciding the placements of these instances. Like in traditional cloud scheduling and cluster scheduling, serverless schedulers need to determine how resources are assigned to user instances in the system based on resource availability. Additionally, serverless schedulers need to make decisions on how to dispatch serverless function invocations to instances. A major thread of research is dedicated to optimizing start time and avoiding unnecessary cold-start latency by optimizing scheduling policies [3, 43, 41]. Besides, as serverless cloud providers are also responsible for dynamically scaling function instances [32], auto-scaling and load-balancing are also important aspects of serverless schedulers. There are research works that consider different kinds of locality [1, 8] to optimize for load balancing performance. Later in this paper, we will show that from an attacker’s view, these are features that can be exploited to increase the co-location attack success rate.
@@ II-B Cloud Co-Location Attacks
Step 3 is the focus of most side-channel attack works, while Steps 0 - 2 are usually neglected by researchers. However, achieving co-location is a non-trivial process, especially as the size of datacenters increases over time. Prior works in recent years [30, 18, 47] indicate that special strategies should be constructed according to scheduler features to tamper with targeted cloud systems.
@@ II-D Threat Model
In this paper, we only study Step 0 (Section III) - Step 1 (Section V) of the co-location attack. We assume non-privileged attackers, i.e., attackers can only launch attacks by invoking functions and collecting the returned execution results. The goal of attackers is to place at least one of their attack function hosts together with a victim’s function host on the same physical machine to enable subsequent attack efforts. In addition, we assume that the attacker does not have any knowledge about the targeted serverless scheduler. Scheduling strategies need to be reverse-engineered (using our method in Section III), based on which the attacker constructs an attack strategy accordingly to target a specific scheduler. However, they can have certain knowledge about the targeted application, e.g., dependencies, which is realistic in the real world. Examples include Airbnb’s disclosure of their usage of open-source projects [5] and Netflix’s publication of their internal tool as an open-source project [36].
@@ III-A Algorithm Description
In this part, we assume a reliable server identification method is provided, i.e., when invoking functions, the attacker will be able to differentiate servers where the function host is placed. This can be based on methods such as time and frequency probing [47] or network-based probing [40]. With this server fingerprinting method, an attacker can continuously ‘discover’ servers in the cloud system and provide unique identifiers to each discovered physical server. Following the same function invocation sequence, the collected trace of such identifiers will be different for different scheduling algorithms, hence they can be used to identify scheduling policies utilized by the cloud provider.
We denote the aforementioned server identification method as a function $\mathcal{F}$. Every time a serverless function $\phi$ is executed in the cluster, a hypothetical fingerprint of the corresponding physical server will be reported [denoted as $\mathcal{F}(\phi)$]. We maintain a mapping record ($\mathcal{R}$) between physical server fingerprints and physical server IDs. When a fingerprint that does not map to any server is returned, a new server in the cluster is discovered. A new ID will be assigned to the server, and a new entry in $\mathcal{R}$ will be established to record the mapping between the server fingerprint and the assigned server ID. The trace of server IDs associated with returned server fingerprints after each function invocation is recorded and will later be used for analysis. As an example, consider a function that is invoked $6$ times. If the first $3$ executions are handled by one server in the cluster, and the last $3$ executions are on another server, the resulting sequence will be $(1,1,1,2,2,2)$.
Shown in Fig. 2, the submission of function invocation consists of $3$ phases (a detailed pseudocode description can be found in Appendix A):
@@
Phase 2. In this step, the attacker constructs a second serverless function $\phi_{1}$, which is a copy of the previous serverless function $\phi_{0}$. Then, the attacker will invoke the two functions alternately for a specific length of time (e.g., 10 minutes), and then wait for all function hosts to be terminated.
@@
Analysis. Features of the scheduling algorithm can be revealed by analyzing and comparing different parts of the collected trace. In Phase 1, the attacker invokes the same function with a relatively high frequency for a sufficiently long pre-defined length of time. The placement sequence in the collected trace can reveal the following features ($\mathrm{F}_{i}$) of the scheduler:
@@
In Phase 2, after all function hosts are terminated, the attacker re-invokes the same function in Phase 1, together with another function with the same functionality but disguised as a different function. This enables the attacker to compare placement results between (1) invocations to the same function that are separated in time, and (2) invocations to different functions. The attacker will be able to answer the following questions:
@@
    Cold-Start Locations: By comparing the placement results of the first function in Phase 1 and Phase 2, the attacker will be able to obtain information regarding the scheduler’s behavior when rescheduling a function instance that has been terminated before. The attacker will be able to answer the following question: for the same function, is the cold-start location always the same?
@@
    Account Locality: By comparing placement results of different functions, the attacker will be able to infer if user locality is considered by the scheduler and answer this question: do the schedulers tend to place function hosts from the same user on the same set of physical servers? This can potentially infect the construction of an attack strategy. This also needs to be cross-checked with placement results from Phase 3.
The goal of Phase 3 is to explore whether there are other locality optimizations that rely on specific configurations of serverless functions (e.g., package required) and are employed by the scheduler. By comparing the placement results of different functions, the attacker will be able to identify:
@@ IV-A Simulator Details
We conduct our fingerprinting experiments in a simulator written in C++. The usage of this simulator allows us to launch large-scale Monte Carlo experiments. The simulator consists of $\sim$2000 Line of Code and simulates the functionality of a serverless system, with a focus on the scheduler. There are $4$ different components in the simulator:
@@
At the beginning of the simulation, a list of users will be randomly generated, and an attacker will be initiated to target a specific user. Both non-malicious users and attackers will generate Function instances when they are initialized. The simulation will go through a pre-defined number of rounds of Function submissions by different users, and the log files will be collected for further analysis. We have implemented different scheduling algorithms in different Scheduler modules.
@@ IV-B Selected Scheduling Algorithms
We selected to implement the following schedulers in both the simulator and the Dask-based serverless platform:
@@
    Helper, which mimics the critical load-balancing behavior of Google Cloud Run [20] described in [47]. For each function to be scheduled, the scheduler selects a server as the base server first. As the load increases, the scheduler creates function host instances on other servers to perform auto-scaling.
@@
    Package-Aware Scheduler (PASch) [8], which is a serverless algorithm that utilizes package locality to improve the performance of serverless platforms. The scheduler maintains a consistent hash ring [24] during operation and maps serverless functions to servers based on the largest package that is required to increase the package cache hit rate in the system and at the same time balance the load of each server in the system.
@@
We utilize OpenWhisk [37] as an example of cold-start-optimized schedulers. Also, as it is one of the most popular open-source serverless platforms, our attack results will be representative. We selected Helper [47] since it is reported to be part of the scheduling algorithm of Google Cloud Run, and we also observe similar behaviors in Microsoft Azure (see Section VII). PASch is selected to represent schedulers with $\mathrm{F}_{5}$ (configuration-based locality) optimizations, and the attack methods on PASch can be easily expanded to target other schedulers with similar optimization features. Finally, Random is selected as the baseline and helps demonstrate if increased randomness can help defend against co-location attacks, as shown by [17] in serverful clouds.
@@ IV-B1 Results
Helper Scheduler. By examining the placement trace in Phase 1, we can clearly observe that the scheduler has optimizations to reduce cold-starts, as functions are invoked repeatedly on the same set of servers ($\mathrm{F}_{1}$). As the number of function invocations increases, the scheduler starts to scale out function hosts by creating more instances to be placed on other servers ($\mathrm{F}_{2}$). In Phase 2, we can obtain the same observation as we start to involve a second function. By comparing the placement results of the overlapped function in Phase 1 and Phase 2, we find that if all hosts of a function are terminated, when re-launching these function hosts, the target scheduler will still choose the same set of servers, i.e., the Helper scheduler presents cold-start locality ($\mathrm{F}_{3}$). Also, we notice that the two different functions in Phase 2 are placed in different locations, indicating the Helper function does not confine the host placement to a small set of servers even though they are from the same user ($\mathrm{F}_{4}$). However, the behaviors of the Helper scheduler in Phase 3 are similar to Phase 1 & 2, indicating that when making scheduling decisions, the scheduler does not consider package locality ($\mathrm{F}_{5}$).
@@
OpenWhisk Scheduler. In Phase 1, when there is only one function involved, we can see that the function executions consistently stay on one machine, indicating that there is optimization to reduce cold-start ($\mathrm{F}_{1}$), yet no auto-scaling is involved ($\mathrm{F}_{2}$). In Phase 2, as we involve a second function, we can see that the number of covered servers increases by 1, indicating that function hosts from the same user are not bonded together ($\mathrm{F}_{4}$), and the location is the same when a function host experiences termination and cold-start ($\mathrm{F}_{3}$). In Phase 3, however, we do not notice the scheduling results change, indicating that the OpenWhisk scheduler does not use package information when making scheduling decisions.
@@
PASch Scheduler. Similar to the OpenWhisk scheduler, in Phase 1, traces collected from PASch show instance locality ($\mathrm{F}_{1}$). However, since the placement consistently stays at the same machine, we can conclude that it does not perform auto-scaling based on the function invocation load ($\mathrm{F}_{2}$). In Phase 2, when a second function is involved, there is still only one server covered. This could potentially be due to (1) user locality, i.e., the scheduler places all functions from a user to a single machine; or (2) the identical configurations of the two functions result in the same placement decision. By examining Phase 3 of the trace, we can exclude possibility (1) since there are different function hosts in this phase. Besides, by comparing scheduling results in Phase 3, an attacker would be able to notice that the scheduler places function hosts with the same package requirements at the same node, demonstrating the existence of package locality-related policies in the scheduling algorithm.